# Chain-of-Influence: Tracing Interdependencies Across Time and Features in Clinical Predictive Modelings

### v1

@@ Abstract
Modeling clinical time-series data is hampered by the challenge of capturing latent, time-varying dependencies among features. State-of-the-art approaches often rely on black-box mechanisms or simple aggregation, failing to explicitly model how the influence of one clinical variable propagates through others over time. We propose Chain-of-Influence (CoI), an interpretable deep learning framework that constructs an explicit, time-unfolded graph of feature interactions. CoI leverages a multi-level attention architecture: first, a temporal attention layer identifies critical time points in a patient’s record; second, a cross-feature attention layer models the directed influence from features at these time points to subsequent features. This design enables the tracing of influence pathways, providing a granular audit trail that shows how any feature at any time contributes to the final prediction, both directly and through its influence on other variables. We evaluate CoI on mortality and disease progression tasks using the MIMIC-IV dataset and a private chronic kidney disease cohort. Our framework significantly outperforms existing methods in predictive accuracy. More importantly, through case studies, we show that CoI can uncover clinically meaningful, patient-specific patterns of disease progression that are opaque to other models, offering unprecedented transparency into the temporal and cross-feature dependencies that inform clinical decision-making.
@@
Chronic Kidney Disease (CKD) affects approximately 8%-16% of the global population and represents one of the most significant challenges in modern healthcare, with its gradual and irreversible progression toward End-Stage Renal Disease (ESRD) imposing substantial clinical and economic burdens National Kidney Foundation (2024); Murphy et al. (2019). The complexity of CKD progression stems from multifaceted interactions between clinical, demographic, and socioeconomic factors that evolve over extended time periods, making accurate prediction and effective management particularly challenging Lin et al. (2013); Li & Padman (2024).
Healthcare data capturing this progression is inherently temporal and multi-dimensional, with clinical outcomes resulting from complex interactions between patient characteristics, laboratory values, medications, and comorbidities. However, accurate prediction alone is insufficient for clinical deployment—clinicians require transparent and interpretable models to confidently incorporate predictive insights into patient care Caruana et al. (2015); Rudin (2019); Tonekaboni et al. (2019). This need for interpretability has driven significant evolution in explainable artificial intelligence (XAI), from traditional post-hoc methods like LIME Ribeiro et al. (2016) and SHAP Lundberg & Lee (2017) toward more sophisticated attention-based approaches that provide intrinsic interpretability Choi et al. (2016; 2017); Bardhan (2024).
Despite these advances, current approaches to clinical prediction face fundamental limitations. Traditional methods either aggregate temporal information into static features, losing crucial temporal dynamics, or model temporal patterns while treating features independently, missing critical inter-feature relationships. Even sophisticated attention-based models like RETAIN Choi et al. (2016), while providing interpretability through temporal and feature-level attention, fail to explicitly model how features influence each other across different time points—a critical gap in understanding disease progression pathways.
@@
Building on TFCAM, we introduce Chain-of-Influence (CoI), which adds hierarchical attention—temporal, feature-level, and cross-feature—and validates these choices with expanded technical details and broader experiments (Li et al., 2025b), enabling principled modeling and interpretation of temporal–feature interdependencies.
@@ 2 Related Work
2.1 Foundational Attention Architectures
@@
Attention mechanisms marked a turning point for interpretability in clinical deep learning. The foundational model in this area, RETAIN, introduced a two-level attention mechanism that operates in reverse chronological order to mimic clinical reasoning Choi et al. (2016). By first weighting the importance of entire patient visits and then individual clinical codes within those visits, RETAIN provides clear, intuitive explanations for its predictions. Following this, models like Dipole extended the concept by using bidirectional LSTMs to capture a more holistic view of the patient’s history, exploring various attention formulations to better understand inter-visit dependencies Ma et al. (2017). While these pioneering models successfully identify what clinical events are important and when they occurred, they model these events largely as independent contributors to the final prediction. They do not explicitly capture the influence pathways—how a variable at one point in time may directly affect a different variable later on.
2.2 Time-Aware and Hierarchical Attention
@@
A significant challenge in clinical data is its inherent irregularity and the varying importance of events over time. To address this, a second wave of research developed more sophisticated temporal attention models. For instance, ATTAIN integrated time-decay factors into its attention weights, allowing the model to dynamically prioritize events based on their temporal proximity to the prediction time Zhang et al. (2019). Other models introduced domain-specific structural priors. StageNet, for example, models disease progression through distinct stages, using a combination of recurrent networks and attention to learn stage-specific representations and their transitions over time Gao et al. (2020). Similarly, AdaCare introduces an adaptive attention mechanism that modulates feature importance based on the patient’s evolving disease severity, captured by a separate deep learning component Ma et al. (2020). These approaches refine the temporal and contextual aspects of attention, but the inter-feature dependencies remain encapsulated within the black-box hidden states of recurrent networks, rather than being an explicit and interpretable output of the model.
2.3 Self-Attention and Transformer-Based Approaches
@@
The advent of the Transformer architecture, with its core self-attention mechanism, provided a powerful tool for capturing global, all-to-all interactions within a sequence Vaswani et al. (2017). This was quickly adapted to the clinical domain in models like BEHRT, which treated a patient’s record as a sequence of concepts and learned deep, contextualized representations Li et al. (2020). Subsequent work built hierarchical structures on this foundation. PAVE, for example, uses multi-layer self-attention to first identify significant clinical events and then aggregate them into higher-level, clinically meaningful event patterns Steinberg et al. (2021). More recently, XTSFormer introduced multi-scale temporal attention to distinguish between the influence of acute, short-term events and chronic, long-term patterns Zhang et al. (2024).
@@
Self-attention excels at modeling a dense web of pairwise interactions. However, this strength can also be a weakness for interpretability in a clinical context. The resulting attention maps highlight a complex network of correlations but do not naturally expose the directed, sequential pathways of influence that characterize disease progression. They show which features co-occur or are correlated, but they do not provide a clear, traceable audit trail of how one event propagates its influence through other variables over time. Our proposed model, CoI, is designed specifically to bridge this gap, moving beyond identifying salient features to explicitly modeling and tracing these inter-feature, cross-time influence chains.
@@ 3 Methodology
In this section, we introduce the Chain-of-Influence (CoI) model, designed to capture and interpret both temporal dynamics and feature interactions in sequential clinical data. CoI integrates three layers of attention mechanisms—temporal attention, feature-level attention, and cross-feature attention—to provide both accurate predictions and rich interpretability.
3.1 Input Representation and Embedding
@@
Given an input tensor $\mathbf{X}\in\mathbb{R}^{B\times T\times F}$ where $B$ is the batch size, $T$ is the number of time steps, and $F$ is the number of input features, the first stage is to project the input into a higher-dimensional embedding space. This is achieved via a learnable linear transformation:
@@
$\mathbf{E}=\operatorname{Linear}(\mathbf{X}),\mathbf{E}\in\mathbb{R}^{B\times T\times D},$
where $D$ denotes the embedding dimension. To encode the temporal order, a fixed positional encoding is added to the embedded inputs, allowing the model to capture sequential information even before applying the attention mechanisms.
3.2 Temporal and Feature-Level Attention
Inspired by the two-level attention mechanism introduced in RETAIN Choi et al. (2016), we employ a dual-attention approach to separately extract temporal and feature-level importance. While RETAIN utilized GRU Cho et al. (2014) units, we implement two distinct bidirectional LSTM branches to better capture long-range dependencies and context from both past and future time steps. This bidirectional processing enables a more comprehensive understanding of temporal patterns, which is particularly crucial for detecting subtle interactions between features across different time points.
Temporal-Level Attention
A bidirectional LSTM is applied to the positionally encoded embeddings: $\mathbf{H}^{\text{temp }}=\operatorname{BiLSTM}(\mathbf{E})$ Subsequently, a linear layer projects the LSTM outputs to compute a scalar attention score $a_{t}$ for each time step: $\alpha_{t}=\operatorname{softmax}\left(\mathbf{w}^{\top}\mathbf{h}_{t}^{\text{temp }}\right)$ These temporal attention weights $\alpha_{t}$ reflect the contribution of each time step to the final prediction.
Feature-Level Attention
In parallel, another bidirectional LSTM processes the same embeddings to capture feature-specific dynamics: $\mathbf{H}^{\text{feat }}=\operatorname{BiLSTM}(\mathbf{E})$ A separate linear projection followed by a non-linear activation (e.g., tanh) is used to compute a vector of feature-level attention weights $\beta_{t}$ for each time step: $\beta_{t}=\tanh\left(\operatorname{Linear}\left(\mathbf{h}_{t}^{\text{feat }}\right)\right)$ Together, the temporal weights $\alpha_{t}$ and the feature-level weights $\beta_{t}$ yield a local contribution matrix $C$, where each element $C[t,i]$ is computed as:
$C[t,i]=\alpha_{t}\times\sum_{k}\left(\beta_{t}[k]\times W_{\mathrm{emb}}[k,i]\times x_{t}[i]\right),$
where $W_{\text{emb }}$ is the weight matrix of the initial embedding layer, and $x_{t}[i]$ is the $i$-th feature of the input at time $t$.
3.3 Cross-Feature Attention and Chained Influence
The cross-feature attention component leverages transformer-based layers to capture temporal dependencies and inter-feature interactions across time. In addition to learning complex patterns, these layers offer a pathway for interpretability by linking local contributions with cross-time interactions. Here, we detail how the cross-attention matrix $A$ is obtained, how the local contribution matrix $C$ is computed, and how they are combined into a chained influence measure.
Deriving the Cross-Attention Matrix AA
Within each transformer layer, multi-head self-attention is applied to the positionally encoded embeddings. For each sample, every layer outputs attention weights with dimensions $\text{num\_heads},T,T]$, where $T$ is the sequence length. These weights indicate the strength of interactions between time steps. To obtain a robust measure of these temporal relationships, we aggregate the attention weights by averaging over all heads and layers, resulting in the cross-attention matrix:
$A\in\mathbb{R}^{T\times T}.$
In our notation, each element $A[t,t^{\prime}]$ represents the attention weight from time step $t^{\prime}$ to time step $t$ (with $t<t^{\prime}$), quantifying how much the information at time step $t$ contributes to the representation at time step $t^{\prime}$.
The matrix $A$ is computed by aggregating the attention weights from all transformer layers and heads. In particular, if $A^{(l,h)}$ denotes the attention weight matrix from the $h$-th head in the $l$-th layer, then we define:
$A[t,t^{\prime}]=\frac{1}{L\cdot H}\sum_{l=1}^{L}\sum_{h=1}^{H}A^{(l,h)}[t,t^{\prime}],$
where $L$ is the number of transformer layers and $H$ is the number of attention heads per layer. This aggregated matrix $A$ captures how information is transmitted from earlier to later time steps. Specifically, a higher value of $A[t,t^{\prime}]$ indicates that the representation at time $t^{\prime}$ heavily incorporates information from time $t$, thereby revealing long-range dependencies and offering insights into the model’s internal temporal dynamics.
Computing the Local Contribution Matrix $C$
@@ 3.3 Cross-Feature Attention and Chained Influence
As detailed in Section 2.1.2, the local contribution matrix $C$ encapsulates the importance of each feature at every time step by integrating the temporal attention weights ($\alpha_{t}$) and the feature-level attention vectors ($\mathbf{b}_{t}$). Combined with the embedding layer’s weight matrix $W_{\mathrm{emb}}$ and the raw input $x_{t}$, the contribution of feature $i$ at time $t$ is computed as:
$C[t,i]=\alpha_{t}\times\sum_{k}\left(b_{t}[k]\times W_{\mathrm{emb}}[k,i]\times x_{t}[i]\right)$
In this expression, the scalar $\alpha_{t}$ captures the overall relevance of time step $t$, while the vector $\mathbf{b}_{t}$ distributes attention across the features at that time. The summation over $k$ aggregates the contribution across the embedding dimensions, effectively translating the feature-level weights into the original input space. Thus, $C[t,i]$ serves as a composite measure that reflects not only the direct impact of feature $i$ at time $t$ but also the modulation of this impact by the importance of the time step and the relative emphasis on the specific feature. It provides a fine-grained interpretation of the direct and modulated impact of feature $i$ at time $t$.
Quantifying Temporal-Feature Cross Influence
Having established the local contribution matrix $C$ and the temporal connectivity matrix $A$, we then seek to quantify explicitly how the importance of a specific feature at an earlier time step affects another feature at a subsequent time step. Specifically, our goal is to detect and measure how feature $i$ at time $t$ influences feature $j$ at time $t^{\prime}$(with $t<t^{\prime}$).

### v2

@@ Abstract
Modeling clinical time-series data is hampered by the challenge of capturing latent, time-varying dependencies among features. State-of-the-art approaches often rely on black-box mechanisms or simple aggregation, failing to explicitly model how the influence of one clinical variable propagates through others over time. We propose Chain-of-Influence (CoI), an interpretable deep learning framework that constructs an explicit, time-unfolded graph of feature interactions. CoI enables the tracing of influence pathways, providing a granular audit trail that shows how any feature at any time contributes to the final prediction, both directly and through its influence on other variables. We evaluate CoI on mortality and disease progression tasks using the MIMIC-IV dataset and a chronic kidney disease cohort. Our framework achieves state-of-the-art predictive performance (AUROC of 0.960 on CKD progression and 0.950 on ICU mortality), with deletion-based sensitivity analyses confirming that CoI’s learned attributions faithfully reflect its decision process. Through case studies, we demonstrate that CoI uncovers clinically meaningful, patient-specific patterns of disease progression, offering enhanced transparency into the temporal and cross-feature dependencies that inform clinical decision-making.
@@
Chronic Kidney Disease (CKD) affects approximately 8%-16% of the global population and represents one of the most significant challenges in modern healthcare, with its gradual and irreversible progression toward End-Stage Renal Disease (ESRD) imposing substantial clinical and economic burdens NKF_CKD_2024; NCHS2019Mortality. The complexity of CKD progression stems from multifaceted interactions between clinical, demographic, and socioeconomic factors that evolve over extended time periods, making accurate prediction and effective management particularly challenging lin2013progression; li2024enhancing.
Healthcare data capturing this progression is inherently temporal and multi-dimensional, with clinical outcomes resulting from complex interactions between patient characteristics, laboratory values, medications, and comorbidities. However, accurate prediction alone is insufficient for clinical deployment—clinicians require transparent and interpretable models to confidently incorporate predictive insights into patient care caruana2015intelligible; rudin2019stop; tonekaboni2019clinicians. This need for interpretability has driven significant evolution in explainable artificial intelligence (XAI), from traditional post-hoc methods like LIME ribeiro2016should and SHAP lundberg2017unified toward more sophisticated attention-based approaches that provide intrinsic interpretability choi2016retain; choi2017gram; bardhan2024icu.
Despite these advances, current approaches to clinical prediction face fundamental limitations. Traditional methods either aggregate temporal information into static features, losing crucial temporal dynamics, or model temporal patterns while treating features independently, missing critical inter-feature relationships. Even sophisticated attention-based models like RETAIN choi2016retain, while providing interpretability through temporal and feature-level attention, fail to explicitly model how features influence each other across different time points—a critical gap in understanding disease progression pathways.
@@
We introduce Chain-of-Influence (CoI), a novel deep learning framework designed to model and interpret these complex temporal-feature interdependencies.
@@ 2 Related Work
Foundational Attention Architectures
@@
Attention mechanisms marked a turning point for interpretability in clinical deep learning. The foundational model in this area, RETAIN, introduced a two-level attention mechanism that operates in reverse chronological order to mimic clinical reasoning choi2016retain. By first weighting the importance of entire patient visits and then individual clinical codes within those visits, RETAIN provides clear, intuitive explanations for its predictions. Following this, models like Dipole extended the concept by using bidirectional LSTMs to capture a more holistic view of the patient’s history, exploring various attention formulations to better understand inter-visit dependencies ma2017dipole. While these pioneering models successfully identify what clinical events are important and when they occurred, they model these events largely as independent contributors to the final prediction. They do not explicitly capture the influence pathways—how a variable at one point in time may directly affect a different variable later on.
Time-Aware and Hierarchical Attention
@@
A significant challenge in clinical data is its inherent irregularity and the varying importance of events over time. To address this, a second wave of research developed more sophisticated temporal attention models. For instance, ATTAIN integrated time-decay factors into its attention weights, allowing the model to dynamically prioritize events based on their temporal proximity to the prediction time zhang2019attain. Other models introduced domain-specific structural priors. StageNet, for example, models disease progression through distinct stages, using a combination of recurrent networks and attention to learn stage-specific representations and their transitions over time gao2020stagenet. Similarly, AdaCare introduces an adaptive attention mechanism that modulates feature importance based on the patient’s evolving disease severity, captured by a separate deep learning component ma2020adacare. These approaches refine the temporal and contextual aspects of attention, but the inter-feature dependencies remain encapsulated within the black-box hidden states of recurrent networks, rather than being an explicit and interpretable output of the model.
Self-Attention and Transformer-Based Approaches
@@
The Transformer and its self-attention mechanism provide a powerful way to model global, all-to-all interactions within a sequence vaswani2017attention. In the clinical domain, BEHRT treats a patient record as a sequence of concepts and learns contextualized representations li2020behrt; PAVE builds on this with multi-layer self-attention to identify important events and aggregate them into higher-level clinical patterns steinberg2021language; and XTSFormer introduces multi-scale temporal attention to separate short-term acute signals from long-term chronic trends zhang2024xtsformer.
@@
While self-attention captures a dense web of pairwise dependencies, the resulting maps remain difficult to interpret clinically: they highlight correlations but do not clearly expose directed, sequential pathways of influence. In particular, they do not provide a traceable audit trail of how an early event propagates through other variables over time. CoI aims to bridge this gap by explicitly modeling and tracing inter-feature influences across time, moving beyond identifying salient visits or features to quantifying their dynamic interactions.
@@ 3 Methodology
In this section, we introduce the Chain-of-Influence (CoI) model, an interpretable deep learning architecture designed to capture and quantify both temporal dynamics and feature interactions in sequential clinical data. The model integrates three complementary attention mechanisms—temporal attention, feature-level attention, and cross-feature attention—into a unified computational graph that enables accurate predictions while providing rich interpretability through explicit influence quantification. Figure 1 provides a complete overview of the model architecture.
Figure 1: CoI architecture overview.
3.1 Model Architecture
3.1.1 Input Representation and Embedding
@@
Given an input tensor $\mathbf{X}\in\mathbb{R}^{B\times T\times D_{\text{in}}}$, where $B$ denotes batch size, $T$ represents the number of time steps, and $D_{\text{in}}$ is the number of input features, we first project the input into a higher-dimensional embedding space through a learnable linear transformation:
$\mathbf{E}=\text{Linear}(\mathbf{X}),\quad\mathbf{E}\in\mathbb{R}^{B\times T\times D_{\text{emb}}},$ (1)
where $D_{\text{emb}}$ denotes the embedding dimension. To encode temporal ordering, we incorporate fixed sinusoidal positional encodings $\text{PE}\in\mathbb{R}^{1\times T\times D_{\text{emb}}}$, yielding positionally-encoded embeddings:
$\mathbf{E}_{\text{t}}=\mathbf{E}+\text{PE},\quad\mathbf{E}_{\text{t}}\in\mathbb{R}^{B\times T\times D_{\text{emb}}}.$ (2)
These embeddings serve as the shared input to three parallel computational pathways described below.
3.1.2 Temporal and Feature-Level Attention
Inspired by the two-level attention framework in RETAIN choi2016retain, we employ two separate bidirectional LSTM modules that operate in parallel on $\mathbf{E}_{\text{t}}$ to extract complementary temporal and feature-level importance scores. Unlike RETAIN’s use of GRU units, our bidirectional LSTMs capture long-range dependencies from both past and future contexts, which is critical for detecting subtle temporal-feature interactions in clinical sequences.
Temporal Attention Module
The first BiLSTM module processes the positionally encoded embeddings to capture temporal dynamics:
$\mathbf{H}^{\text{temp}}=\text{BiLSTM}_{\alpha}(\mathbf{E}_{\text{t}}),\quad\mathbf{H}^{\text{temp}}\in\mathbb{R}^{B\times T\times 2H},$ (3)
where $H$ is the hidden dimension of each LSTM direction. The concatenated forward and backward hidden states are then projected to scalar attention scores via a linear layer followed by sigmoid activation:
@@ 3.1.2 Temporal and Feature-Level Attention
$\bm{\alpha}=\sigma(\text{Linear}_{2H\to 1}(\mathbf{H}^{\text{temp}})),\quad\bm{\alpha}\in\mathbb{R}^{B\times T\times 1},$ (4)
where $\sigma(\cdot)$ denotes the activation function. Each element $\alpha_{t}\in(0,1)$ quantifies the importance of time step $t$ to the final prediction. Critically, $\bm{\alpha}$ is computed during the forward pass but applied later (see Equation 10), enabling end-to-end gradient flow.
Feature Attention Module
In parallel, a second BiLSTM branch extracts feature-specific temporal patterns:
$\mathbf{H}^{\text{feat}}=\text{BiLSTM}_{\beta}(\mathbf{E}_{\text{t}}),\quad\mathbf{H}^{\text{feat}}\in\mathbb{R}^{B\times T\times 2H}.$ (5)
A separate linear projection followed by sigmoid activation produces feature-level attention vectors:
$\bm{\beta}=\sigma(\text{Linear}_{2H\to D_{\text{emb}}}(\mathbf{H}^{\text{feat}})),\quad\bm{\beta}\in\mathbb{R}^{B\times T\times D_{\text{emb}}},$ (6)
where each $\beta_{t}[d]\in(0,1)$ represents the importance of embedding dimension $d$ at time $t$. Unlike temporal attention, feature attention is applied immediately to modulate the embeddings before they enter the transformer layers.
The feature attention weights $\bm{\beta}$ are applied to $\mathbf{E}_{\text{t}}$ via element-wise multiplication:
$\widetilde{\mathbf{E}}=\mathbf{E}_{\text{t}}\odot\bm{\beta},\quad\widetilde{\mathbf{E}}\in\mathbb{R}^{B\times T\times D_{\text{emb}}},$ (7)
where $\odot$ denotes element-wise multiplication. This operation allows the model to selectively emphasize relevant feature dimensions before cross-feature interactions are learned in subsequent transformer layers.
3.1.3 Cross-Feature Transformer Layers
The feature-weighted embeddings $\widetilde{\mathbf{E}}$ are passed through $L$ stacked Transformer layers to model temporal dependencies and inter-feature interactions. Each layer $l$ applies multi-head self-attention ($N_{h}$ heads) and a feed-forward network, with residual connections and Dynamic Tanh (DyT) normalization zhu2025transformers after each sub-layer.
Given input $\mathbf{Z}^{(l-1)}\in\mathbb{R}^{B\times T\times D_{\text{emb}}}$ (with $\mathbf{Z}^{(0)}=\widetilde{\mathbf{E}}$), the attention output is $\mathbf{U}^{(l)}$, and attention weights from all layers are retained for interpretability (Section 3.2). The attention sub-layer is
$\mathbf{Z}_{\text{attn}}^{(l)}=\text{DyT}\left(\mathbf{Z}^{(l-1)}+\text{Dropout}(\mathbf{U}^{(l)})\right),$ (8)
where $\text{DyT}(\mathbf{x})=\alpha\cdot\tanh(\beta\cdot\mathbf{x}+\gamma)$ with learnable $\alpha,\beta,\gamma$. The feed-forward sub-layer uses a two-layer MLP with ReLU and hidden size $4D_{\text{emb}}$:
$\mathbf{Z}^{(l)}=\text{DyT}\left(\mathbf{Z}_{\text{attn}}^{(l)}+\text{Dropout}(\text{FFN}(\mathbf{Z}_{\text{attn}}^{(l)}))\right).$ (9)
After $L$ layers, $\mathbf{Z}^{(L)}\in\mathbb{R}^{B\times T\times D_{\text{emb}}}$ encodes rich temporal–feature interactions.
3.1.4 Temporal Attention, Pooling, and Output Projection
The temporal attention weights $\bm{\alpha}$ computed in Equation 4 are applied to the transformer output via element-wise multiplication, then aggregated via global average pooling, and finally passed through MLPs to produce predictions:
$\mathbf{\tilde{Z}}$ $=\mathbf{Z}^{(L)}\odot\bm{\alpha},\quad\mathbf{\tilde{Z}}\in\mathbb{R}^{B\times T\times D_{\text{emb}}},$ (10)
$\mathbf{z}$ $=\frac{1}{T}\sum_{t=1}^{T}\mathbf{\tilde{Z}}[:,t,:],\quad\mathbf{z}\in\mathbb{R}^{B\times D_{\text{emb}}},$ (11)
$\hat{\mathbf{y}}$ $=\text{MLP}(\mathbf{z}),\quad\hat{\mathbf{y}}\in\mathbb{R}^{B\times D_{\text{out}}},$ (12)
where MLP is a two-layer feed-forward network with ReLU activation.
3.2 Interpretability via Chain-of-Influence
Beyond accurate predictions, the CoI model provides rich interpretability by explicitly quantifying how features at earlier time steps influence features at later time steps. This is achieved by combining the local contribution matrix $\mathbf{C}$ (derived from dual attention mechanisms) with the cross-attention matrix $\mathbf{A}$.
3.2.1 Local Contribution Matrix
@@
The local contribution matrix $\mathbf{C}\in\mathbb{R}^{B\times T\times D_{\text{in}}}$ quantifies the importance of each input feature at every time step by integrating temporal attention weights $\bm{\alpha}$, feature-level attention vectors $\bm{\beta}$, and the embedding layer’s weight matrix $\mathbf{W}_{\text{emb}}$. For each sample $n$, time step $t$, and input feature $i$:
$C[n,t,i]=\alpha[n,t]\times\sum_{k=1}^{D_{\text{emb}}}\left(\beta[n,t,k]\times W_{\text{emb}}[k,i]\times X[n,t,i]\right),$ (13)
where the summation over embedding dimensions $k$ aggregates the contribution back to the original input space. This formulation captures both the temporal importance (via $\alpha_{t}$) and feature-specific emphasis (via $\beta_{t}$), providing a fine-grained interpretation of direct feature impacts.
3.2.2 Cross-Attention Matrix
To quantify how information propagates across time steps, we aggregate the attention weights from all transformer layers and heads. For each sample, layer $l$, and head $h$, the attention weights form a matrix $\mathbf{A}^{(l,h)}\in\mathbb{R}^{T\times T}$, where $A^{(l,h)}[t,t^{\prime}]$ represents the attention from time step $t^{\prime}$ to time step $t$. We compute the averaged cross-attention matrix as:
$A[t,t^{\prime}]=\frac{1}{L\cdot N_{h}}\sum_{l=1}^{L}\sum_{h=1}^{N_{h}}A^{(l,h)}[t,t^{\prime}],$ (14)
where higher values of $A[t,t^{\prime}]$ indicate stronger temporal dependencies from time $t^{\prime}$ to time $t$.
3.2.3 Chain of Influence Measure